### v1

# Physics motivated models of pulsar X-ray hotspots: off-center dipole configurations

Abstract
Recently, it was proposed that an off-center dipole magnetic configuration, together with a non-trivial temperature profile, may be the best model to explain the X-ray light curve of PSR J0030+0451 observed by the Neutron Star Interior Composition Explorer (NICER). Using a theoretical model for the electric current density in a force-free pulsar magnetosphere, we compute from first principles the distribution of electric current over the polar cap associated with an off-center magnetic dipole. We then use a simple prescription to compute the resulting temperature distribution, which allows us to derive the observed X-ray light curve. We investigate the role of the volumetric return current region in the polar cap and find that although it does not make a big difference in an aligned dipole case, the difference can be bigger in the case of an off-center dipole. Finally, we apply Markov Chain Monte Carlo (MCMC) fitting to the X-ray light curves of pulsars PSR J0030+0451 and PSR J0437–4715 with and without the volumetric return current, and find that our model can reasonably recover the observed X-ray light curves.
1 Introduction
X-ray pulse profiles modeling constitutes a crucial methodology for deducing key parameters of neutron stars such as their mass and radius. Typically, this type of modeling involves reverse-engineering observed X-ray light curves and spectra to infer the temperature distribution on the stellar surface, e.g. the locations and temperatures of hotspots (see e.g. Morsink et al., 2007; Bauböck et al., 2013; Psaltis & Özel, 2014; Vincent et al., 2018).
With the advent of the Neutron star Interior Composition Explorer (NICER) and its high-quality timing data, precision modeling of the X-ray pulse profiles of millisecond pulsars became a possibility. Miller et al. (2019) and Riley et al. (2019) made the first measurement of the mass and radius of PSR J0030+0451 by applying detailed modeling techniques to NICER data. The mass-radius measurement holds substantial implications, providing a means to constrain the equation of state for matter inside the neutron star, probing nuclear physics at these extreme conditions that are not accessible on Earth (See e.g. Raaijmakers et al., 2019, 2020, 2021; Huang et al., 2024, 2024; Rutherford et al., 2024).
In addition to determining the mass and radius of PSR J0030+0451, Riley et al. (2019) produced, at the time, the most detailed hotspot map of the neutron star surface derived from observational data. The hotspots were found to have distinct shapes: one is approximately circular, while the other is like a crescent. Both of the hotspots were found to be located within the same hemisphere, suggesting magnetic structures beyond the regular dipole close to the star (Bilous et al., 2019). A few theoretical works followed, attempting to study the implication of the hotspot shapes on the magnetic field structure near the star (Chen et al., 2020; Kalapotharakos et al., 2021), and finding that a dipole plus quadrupole magnetic field can typically reproduce the observed X-ray light curve.
A recent investigation by Vinciguerra et al. (2024) revisited the hotspot modeling of PSR J0030+0451, finding the hotspots to be consistent with the polar caps of an off-center dipole magnetic field but with a nontrivial temperature distribution. In particular, the new study found that one of the hotspots has a high-temperature component near its edge. This study suggests that nontrivial temperature distribution across hotspots may need to be taken into account during the modeling process, which may significantly expand the parameter space. A theory that describes the temperature distribution across pulsar surface hotspots is urgently needed.
Fortunately, recent advances in theoretical modeling of the pulsar magnetosphere have provided a basis for such a theory. Gralla et al. (2017) have outlined a model to compute the magnetospheric current distribution across the neutron star surface, and Lockhart et al. (2019) have applied this model to compute potential surface temperature maps given a dipole plus axisymmetric quadrupole magnetic field configuration. However, the analytic models based on magnetospheric current have been limited to setups that are symmetric with respect to some magnetic axis. None of the hotspot configurations from phenomenological X-ray light curve modeling (e.g. Miller et al., 2019; Riley et al., 2019) are axisymmetric. Going beyond axisymmetric models is a crucial step to take in order to bridge the gap between our theoretical understanding of pulsar magnetospheres and the observed X-ray data.
In this paper, we attempt to take a first step towards a non-axisymmetric, physics-based model of the temperature distribution across hotspots on the surfaces of rotation-powered pulsars. Based on the theory of force-free pulsar magnetospheres developed by Gralla et al. (2017), we consider an off-center magnetic dipole and derive the distribution of the current density across the neutron star surface. Using this current distribution, we compute the surface heating rate using a simple prescription and derive the surface temperature distribution. Finally, we compute the pulsar’s light curve using the X-PSI global temperature module and incorporate it into pulse-profile modeling process. We perform a direct fit to the observed X-ray light-curve data of PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437), aiming to showcase the explanatory power of our off-center dipole temperature map. This work illustrates the potential of adopting a physics-based hotspot configuration to model pulsar X-ray data, offering an alternative to the geometric methods currently in use (e.g., those proposed by Riley et al., 2019).
This paper is organized as follows. In Section 2, we provide a detailed description of the theoretical method. In Section 3, we compare the surface current distribution for shifted dipole models with results from canonical centered dipole models, highlighting the differences in hotspot structure and resulting X-ray light curves. In Section 4, we apply this model to try to fit the light curves of millisecond pulsars J0030 and J0437. Finally in Section 5 we discuss the limitations of this method and outline future directions for this line of work.
2 Method
We follow the general method developed by Gralla et al. (2017) and Lockhart et al. (2019) to compute the current distribution on the surface of the pulsar. In this section, we outline the key ingredients, while referring the reader to the original paper for more detail.
The pulsar magnetosphere is assumed to be well-described by force-free electrodynamics (see e.g. Gruzinov, 1999; Contopoulos et al., 1999). Under such an assumption, the electric current flowing in the magnetosphere is entirely determined by the magnetic field configuration. Gralla & Jacobson (2014) showed that in this limit the electromagnetic field can be described using only two scalar fields $\alpha$ and $\beta$, which are often called “Euler potentials” (Uchida, 1997):
$F=d\alpha\wedge d\beta.$ (1)
Magnetic field lines are given by the intersection of surfaces of constant $\alpha$ and $\beta$. The Euler potentials satisfy the simple geometric equation:
$\alpha\wedge d*F=\beta\wedge d*F=0.$ (2)
Under axisymmetry, one of the Euler potentials becomes trivial, $\beta=\phi$, and the other equation can be reduced to the well-known Grad-Shafranov equation or pulsar equation (see e.g. Michel, 1973; Contopoulos et al., 1999) Solving the full non-axisymmetric problem in full 3D is still a challenging problem and often done using time-dependent numerical simulations (e.g. Spitkovsky, 2006).
If a solution is already found, Gralla et al. (2017) showed that the four-current density in the pulsar magnetosphere can be computed directly from the Euler potentials:
$J^{\hat{t}}$ $=\sqrt{1-\frac{2M}{r}}J^{t},$ (3)
$J^{\hat{r}}$ $=\frac{\Lambda(\alpha,\beta)\left(\partial_{\theta}\alpha\partial_{\phi}\beta-\partial_{\phi}\alpha\partial_{\theta}\beta\right)}{\sqrt{r(r-2M)}(r\sin\theta)},$ (4)
$J^{\hat{\theta}}$ $=\frac{\Lambda(\alpha,\beta)}{r\sin\theta}\left(\partial_{\phi}\alpha\partial_{r}\beta-\partial_{r}\alpha\partial_{\phi}\beta\right),$ (5)
$J^{\hat{\phi}}$ $=\frac{\Lambda(\alpha,\beta)}{r}\left(\partial_{r}\alpha\partial_{\theta}\beta-\partial_{\theta}\alpha\partial_{r}\beta\right),$ (6)
Where $(J^{\hat{t}},J^{\hat{r}},J^{\hat{\theta}},J^{\hat{\phi}})$ are the components of the four-current components under the orthonormal basis $(\hat{\bm{t}},\hat{\bm{r}},\hat{\bm{\theta}},\hat{\bm{\phi}})$, and $J^{t}$ is given by:
$\begin{split}J^{t}=&\frac{\Omega-\Omega_{Z}}{r(r-2M)}\left\{\partial_{\theta}\alpha\partial_{\theta}\partial_{\phi}\beta-\partial_{\theta}\beta\partial_{\theta}\partial_{\phi}\alpha\right.\\ &+r(r-2M)\left(\partial_{r}\alpha\partial_{r}\partial_{\phi}\beta-\partial_{r}\beta\partial_{r}\partial_{\phi}\alpha\right)\\ &-\partial_{\phi}\alpha\left[\left(1-\frac{2M}{r}\right)\partial_{r}\left(r^{2}\partial_{r}\beta\right)\right.\\ &\left.+\frac{\partial_{\theta}\left(\sin\theta\partial_{\theta}\beta\right)}{\sin\theta}\right]+\partial_{\phi}\beta\left[\left(1-\frac{2M}{r}\right)\right.\\ &\left.\partial_{r}\left(r^{2}\partial_{r}\alpha\right)+\frac{\partial_{\theta}\left(\sin\theta\partial_{\theta}\alpha\right)}{\sin\theta}\right]\},\end{split}$ (7)
where $\Omega$ is the angular velocity, $\Omega_{Z}$ is the “frame-drag frequency” defined as $\Omega_{Z}=2\hat{I}\Omega/r^{3}$, and $\hat{I}$ is the moment of inertia, which defined as $\hat{I}=\mathcal{I}M_{\star}R_{\star}^{2}$, where we set the compactness $\mathcal{I}=2/5$. $M_{\star}$ and $R_{\star}$ are the mass and radius of the neutron star respectively. The function $\Lambda(\alpha,\beta)$ is a conserved quantity along each field line that determines the magnitude of the current density. Comparing with first-principles PIC simulations of the global pulsar magnetosphere, Gralla et al. (2017) were able to find an approximate expression for this function:
$\Lambda(\alpha,\beta)=$ $\mp 2\Omega\left\{J_{0}\left(2\arcsin\sqrt{\alpha/\alpha_{o}}\right)\cos\iota\right.$ (8)
$\left.\mp J_{1}\left(2\arcsin\sqrt{\alpha/\alpha_{o}}\right)\cos\beta\sin\iota\right\},$
$\alpha<\alpha_{o},$
where the $\mp$ sign refers to the northern/southern hemisphere, $\iota$ is the dipole inclination angle, $J_{0}$ and $J_{1}$ are Bessel functions of the first kind, and $\alpha_{0}$ label the last open field line, defined as:
$\alpha_{o}=\sqrt{\frac{3}{2}}\mu\Omega\left(1+\frac{1}{5}\sin^{2}\iota\right).$ (9)
This quantity $\alpha_{0}$ delineates the boundary of the polar cap, which is the region where open magnetic field lines originate from. The dipole magnetic moment $\mu$ is defined as $B_{0}\,R_{\star}^{3}$, where $B_{0}$ represents the magnetic field strength on the surface of the neutron star. In the subsequent sections, we adopt a value of $B_{0}=5\times 10^{8}\,\mathrm{G}$. Equation (9) is our definition of the pulsar polar cap for the rest of this paper, even when the designated regions are no longer geometrically similar to polar caps in the traditional sense.
For a general rotating neutron star, no known analytic solution exists for the force-free Euler potentials $\alpha$ and $\beta$. However, if we only consider a non-relativistic dipole-like magnetic field originating from the star, then a simple analytic solution exists for the Euler potentials:
$\alpha=-\frac{\mu}{r}\sin^{2}\theta^{\prime},\quad\beta=\varphi^{\prime},$ (10)
where $\theta^{\prime}$ and $\varphi^{\prime}$ are the polar coordinates about the dipole axis, and they are related to the coordinates $\theta$ and $\varphi$ in the lab frame by:
$\cos\theta^{\prime}$ $=\cos\theta\cos\iota-\sin\theta\cos\varphi\sin\iota$ (11)
$\tan\varphi^{\prime}$ $=\frac{\sin\theta\sin\varphi}{\sin\theta\cos\varphi\cos\iota+\cos\theta\sin\iota},$ (12)
where $\iota$ is the inclination angle between the magnetic axis and the rotation axis.
It is straightforward to generalize Equation (10) to a shifted dipole configuration. Define the shifted coordinates as:
$x_{s}=x+x_{0},\quad y_{s}=y+y_{0},\quad z_{s}=z+z_{0},$ (13)
where $(x_{0},y_{0},z_{0})$ is the center of the shifted dipole,$(x,y,z)$ is the Cartesian coordinate with respect to the shifted magnetic center, and $(x_{s},y_{s},z_{s})$ is the Cartesian coordinate with respect to the actual center of the neutron star. Equations (10) are valid in the magnetic body frame $(x^{\prime},y^{\prime},z^{\prime})$, whereas the derivatives in Equations (3)–(7) now need to be taken with respect to the lab frame $(r_{s},\theta_{s},\varphi_{s})$.
Note that we have adopted the expression for a non-relativistic magnetic dipole in Equations (10). Gralla et al. (2016) has presented a general-relativistic solution of the magnetic dipole in terms of Euler potentials. However, once the magnetic dipole is shifted from the center of mass of the neutron star, the general relativistic solution no longer applies. We have verified that even for millisecond pulsars, the GR solution given by Gralla et al. (2016) only differs from the non-relativistic dipole solution by around $3\%$. On the other hand, the frame-dragging effect due to the rotation of the star, which is the dominant effect of general relativity, is already built into the prescription of computing the current density given in Equations (3)–(7).
The body frame coordinates $(r,\theta^{\prime},\varphi^{\prime})$ are related to the lab frame coordinates $(r_{s},\theta_{s},\varphi_{s})$ by a nonlinear transformation. As a result, taking the derivatives of $\alpha$ and $\beta$ with respect to $(r_{s},\theta_{s},\varphi_{s})$ is quite nontrivial. We overcome this difficulty by employing the Automatic Differentiation (AD) technique while computing the Euler potentials. AD keeps a graph of mathematical operations during the calculation of $\alpha$ and $\beta$, then one can easily and accurately take the partial derivatives of these complicated functions with respect to the coordinates by walking down the computational graph and applying the chain rule. Automatic differentiation allows us to achieve machine precision when evaluating the derivatives, without the need to introduce a grid. This is especially important in Section 4 where we need to evaluate the derivatives a large number of times to carry out the Markov Chain Monte Carlo (MCMC) fitting procedure. We use the open source Python automatic differentiation package jax [^1] https://github.com/google/jax to compute the numerical derivatives of the Euler potentials in the production code.
Given the four-current density across the polar cap, a heating prescription is required to compute the temperature distribution. This heating is intimately related to the dissipation of electromagnetic energy, particle acceleration, and the pair production process above the neutron star surface. As shown through analytic calculations by of Beloborodov (2008) and simulations by Timokhin & Arons (2013), particle acceleration at the polar cap depends on the ratio $j/c\rho_{\mathrm{GJ}}$, where $j$ is the magnitude of the three-current density and $\rho_{\mathrm{GJ}}$ is the Goldreich-Julian charge density Goldreich & Julian (1969). If this ratio is locally greater than 1 (the four-current is spacelike, $J^{2}>0$) or less than 0 ($j$ and $\rho_{\mathrm{GJ}}$ have opposite signs), these regions can build up a large voltage, accelerating particles to very high Lorentz factors, triggering $e^{\pm}$ pair production. A fraction of the produced pairs precipate on the polar cap, converting some of the energy dissipated in the gap to heat the neutron star surface. On the other hand, if $10<j/c\rho_{\mathrm{GJ}}<$ 1 (the four-current is timelike, $J^{2}<0$), the plasma can flow along the field line with a mildly relativistic Lorentz factor without building up a large voltage (Chen & Beloborodov, 2013).
In this paper, we follow the simple prescription used by Lockhart et al. (2019), where the surface heating power $P$ is proportional to the magnitude of the three-current density $j$. Equating the heating rate with the radiation power gives a simple relation between current density and surface temperature:
$T(\theta,\phi)\equiv\left\{\begin{array}[]{lll}T_{0}\left|\bm{j}/j_{0}\right|^{1/4},&|\rho|<|\bm{j}|\text{ or }\rho_{\mathrm{GJ}}j_{r}<0\\ 0,&|\rho|>|\bm{j}|\end{array}\right.$ (14)
Here the $j_{0}$ and $T_{0}$ determine the normalization of current density and temperature on neutron star, and the combination $T_{0}/j_{0}^{1/4}$ can be treated as a fitting parameter. In the subsequent sections, we set this constant to unity, since we are only interested in the shape of pulse profile and this constant only affects the normalization of the X-ray flux. As a result, the temperature values computed in this paper are in arbitrary numerical units.
The region where $\rho_{\mathrm{GJ}}j_{r}<0$ is called the (volumetric) “return current” region which always exists in the force-free solution (see e.g. Parfrey et al., 2012), independent of the return current sheet along the separatrix between the open and closed field line zones. Previous modeling work such as Lockhart et al. (2019) mostly focused on the spacelike region where $j/c\rho_{\mathrm{GJ}}>1$, and ignored the return current region where $j$ and $\rho_{\mathrm{GJ}}$ have opposite signs. In this paper, we treat both regions using the same heating model described by Equation (14), in order to asses the importance of modeling the return current region. As can be seen in Sections 3 and 4, the return current region may contribute a non-negligible amount to the X-ray light curves, influencing the results of the fitting process.
Figure 1: Temperature distribution of different hotspost configurations (magnetic inclination $\iota=45^{\circ}$.):(a) Uniform centered dipole circular hotspots with $T=10^{6}\,K$.(b) The centered dipole with computed temperature distribution without considering return current. (c) The centered dipole with computed temperature distribution with considering return current region. All the temperature profile normalized by the maximum temperature of this profile. The cyan boundary shared by a,b,c represent the polar cap boundary delineated by $\alpha=\alpha_{0}$ from Equation (9).
etnec
Figure 2: Pulse profile computed from the centered dipole field in three cases: Uniform hotspot temperature (red), Without return current (blue), With return current (green). All the other parameters to generate this plot are listed in Table 1. The magnetic inclination angle $\iota=45^{\circ}$, observer inclination angle $I=90^{\circ}$.
To compute the X-ray light curves from the surface temperature profiles, we use the open source X-PSI package by Riley et al. (2019) (https://github.com/xpsi-group/xpsi.git, Riley et al. (2023)), which follows a methodology developed over many previous works, including Bogdanov et al. (2019); Riley et al. (2019); Bogdanov et al. (2021); Riley et al. (2021). To describe the shape of the hotspots, the X-PSI package normally utilizes a model that represents the hot-spot surface through overlapping spherical caps with uniform temperature, but it also has the capability to accommodate irregular shapes of hot-spots by specifying the temperature distribution over the whole stellar surface  (see e.g. Das et al., 2024). It is the later mode that we are using in the calculations presented below. For definitiveness, the physical parameters we used for computing the light curves are summarized in Table 1.
In the closed field line zone of the pulsar magnetosphere, no current is flowing along the magnetic field lines. As a result, our simple model would predict there is zero heating in these regions which often occupies a large portion of the stellar surface. For these areas, a uniform temperature of $10^{3}$ in our numerical units is assigned as the background temperature when computing the sample light curve. We keep this surface temperature $T_{s}$ a fitting parameter for Section 4. The choice of fitting parameters will be discussed in more detail in Section 4.1.
3 Results
3.1 Centered Dipole
Figure 3: 4-current density squared $J^{2}$ and temperature $T$ distribution of shifted hotspot projected on the star surface, with magnetic inclination $\iota=45^{\circ}$ and dipole shift $\delta=R_{\star}/2$ along different directions. Panels (a) and (b): $x$–direction shift; panels (c) and (d): $y$–direction; panels (e) and (f): $z$–direction shift. Each color map is normalized by the maximum value of $J^{2}$ and $T$.
Figure 4: Pulse profile computed from the shifted dipole field in three cases: $x$–direction shift (blue), $y$–direction shift (red) and $z$–direction shift (green), comparing with centered dipole field computation without including any shift (black dashed). All the other parameters to generate this plot are listed in Table 1 with magnetic inclination $\iota=45^{\circ}$, observer inclination $I=90^{\circ}$
In order to validate our method, we first compute the polar cap temperature distribution for a centered magnetic dipole. Figure 1 compares the temperature distribution for a centered dipole at inclination angle $\iota=45^{\circ}$, using three different methods for computing the hotspot temperature distribution: uniform temperature of $T=10^{6}\,K$, only spacelike region ($j/c\rho_{\mathrm{GJ}}>1)$, and spacelike region plus volumetric return current ($j/c\rho_{\mathrm{GJ}}<0$). The shapes of the activated polar caps agree with previous studies such as Bai & Spitkovsky (2010); Gralla et al. (2017). We find that the volumetric return current region always produces a small crescent-shaped region with much lower temperature than the main polar cap.
We also computed the pulse profiles of each case assuming an observer at observer inclination $I=90^{\circ}$. The results are shown in Figure 2. Similar to Lockhart et al. (2019), we do not observe significant deviation between all three pulse profiles, The elevated amplitude observed in the light curve derived from the uniform temperature computation is primarily attributable to our specific choice of temperature. Determining an appropriate uniform temperature that accurately represents the underlying physics is challenging; therefore, we have adopted a value of $10^{6}\,K$ solely for demonstrative and comparative purposes. Although the uniform temperature model produces a light curve very similar to our model with more realistic hotspot configuration, it is but a coincidence in our choice of parameters.
In particular, Figure 2 shows that the contribution of the volumetric return current region is quite subdominant, justifying the choice of ignoring this component in previous studies. However, in our subsequent investigations of a shifted dipole magnetic configuration, there is no reason a priori to expect the return current continues to play a subdominant role. Therefore, we will keep this component in all subsequent calculations to assess its importance.
3.2 Shifted Dipole
We now apply the methods described in Section 2 to compute the temperature maps and X-ray light curves of shifted dipole configurations. In this section, our goal is to demonstrate the feasibility of our method, only exploring the impact of dipole displacement on the current and temperature distribution for specific parameters. In particular, we examine a magnetic dipole shifted in $x$, $y$, and $z$ direction respectively, each by a distance of $\delta=R_{*}/2$. This shift is intentionally chosen to be large, in order to show its dramatic effects on both the current density and temperature profiles, as well as on the X-ray pulse profile.
Figure 3 illustrates the effect of a shift $\delta=R_{\star}/2$ along the $x$, $y$, and $z$ directions on the polar cap 4-current density structure and temperature distribution. For example, in the panel (a), we present the $x$–direction shift field 4-current density distribution across the polar cap on the stellar surface. Compared to the centered dipole configuration, the north polar cap is greatly extended to a larger area, with a single hotspot at its center. In contrast, the south polar cap is compressed into a much smaller region. It should be noted that the small blank region at the center of the north pole is an artifact of the projection used in the plot. In reality, the hot spot does not exhibit any hollow feature. The panel (b) shows that the south polar cap , despite being much smaller, has a significantly higher temperature than the north. This is consistent with the expectation that regions located closer to the dipole center experience stronger magnetic fields and thus have stronger heating. Similar asymmetry between the temperatures of the two polar caps can be seen in the $z$–shifted dipole case as well (see panels (e) and (f) of Figure 3).
The four-current density and temperature distribution for the $y$–shift scenario $\delta_{y}=R_{\star}/2$ in a dipole field are shown in the panels (c) and (d) of Figure 3. The two polar caps remain symmetric, and generally similar to the unshifted case shown in Figure 1. The main difference is that the hot spots become closer together compared to the centered dipole scenario. This displacement changes the distance between the two peaks in the X-ray light curve in one rotation period, as can be seen in Figure 4. Such a distinct outcome for a $y$–direction shift arises naturally because the star’s dipole moment lies in the $x-z$ plane, so shifting perpendicular to that plane yields behavior not seen in within-plane shifts.
The panels (e) and (f) of Figure 3 show how a shift $\delta_{z}=R_{\star}/2$ along the $z$–direction affects the hotspot current density distribution and temperature map. In panel (e), the two polar caps display noticeably different areas, with the north polar cap undergoing an enlargement similar to what was observed in the $x$–shift scenario, and southern polar cap also exhibited a larger current density similar with what we see in $x$–direction shift. The panel (f) reveals that, in this configuration, the south hotspot appears significantly brighter, whereas the north hotspot—despite having a larger area— is comparatively fainter, similiar to the features seen in the $x$–shift case.
The X-ray pulse profiles for shifts in the $x$–, $y$–, and $z$–directions are computed and presented in Figure 4 for a neutron star with magnetic inclination angle $\iota=45^{\circ}$, as seen by an observer at angle $I=90^{\circ}$ with respect to the rotation axis. Additional parameters, such as the mass and radius of the neutron star, are required to determine its external spacetime, thereby allowing us to generate the pulse profiles using the X-PSI package. We used the parameters of the pulsar J0030+0451 for these light curves, and the values are summarized in Table 1.
Without any normalization added, our model indicates that for an $x$–direction shift, X-ray emission from the south polar cap dominates, yielding a significantly larger peak-to-valley ratio than the centered-dipole light curve (shown as the dashed curve). Although the south polar cap is smaller in area, it is brighter due to the locally enhanced magnetic field. Consequently, the maximum peak in the $x$–shift configuration is nearly twice that of the centered-dipole setup. In the $y$–direction shift scenario, the north and south polar caps move closer, merging into a single, highly boosted emission peak; Figure 4 shows that this configuration produces both the highest peak and the largest peak-to-valley ratio, implying a higher total flux than in the other shift cases. Finally, the $z$–direction shift presents an interesting contrast. Although the north polar cap appears fainter than the south cap in Figure 3, the observer inclination $I=90^{\circ}$ (equatorial viewing) brings the north cap closer to the line of sight. As a result, the pulse profile exhibits the smallest peak-to-valley ratio, and the south cap’s higher temperature does not substantially exceed the emission level seen in the centered-dipole configuration.
        Model Parameters         PSR J0030+0451         PSR J0437–4715
        Mass $M_{\star}$         1.385 $M_{\odot}$         1.418 $M_{\odot}$
        Radius $R_{\star}$         11.70 km         11.36 km
        Earth distance $D$         0.2 kpc         156.98 pc
        $\nu$         205.3 Hz         174 Hz
        $B$         $5\times 10^{8}$ G         $5\times 10^{8}$ G
Table 1: This is a summary for all the pulse-profile modeling parameters setting for (a) PSR J0030+0451 and (b) PSR J0437–4715. The J0030 parameters are also used in Section 3 to compute the hot spot temperature distribution of centered and shifted dipoles. Mass $M_{\star}$ Radius $R_{\star}$ refers the neutron star mass and radius, earth distance $D$ defined as the distance between this source and earth, $\nu$ is the Coordinate frequency of the mode of radiative asymmetry in the photosphere that is assumed to generate the pulsed signal [Hz]. $B$ in G [Gauss], is the surface magnetic field that take into the computation of magnetic moment.
      Model Parameters       Prior
      observer inclination $I$       $\mathcal{U}(0,\frac{\pi}{2})$
      magnetic inclination $\iota$       $\mathcal{U}(0,\frac{\pi}{2})$
      $x_{0}$       $\mathcal{U}(-R_{\star},R_{\star})$
      $y_{0}$       $\mathcal{U}(-R_{\star},R_{\star})$
      $z_{0}$       $\mathcal{U}(-R_{\star},R_{\star})$
      $\log_{10}T_{s}$       $\mathcal{U}(0,10)$
Table 2: Summary of Inference Model Parameters and Prior Settings: The observer inclination is denoted by $I$ and the magnetic inclination by $\iota$. The off-center dipole displacements along the $x$–, $y$–, and $z$–axes are represented by $x_{0}$, $y_{0}$, and $z_{0}$, respectively. $T_{s}$ denotes the temperature of the neutron star’s non-hotspot region (in arbitrary units). Here, $\mathcal{U}$ denotes a uniform distribution.
Figure 5: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shifts and $\log_{10}T_{s}$ after applying constraints to the observed pulse profile of J0030. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while $T_{s}$ is in numerical units.
Highest Likelihood Case 1 Case 2
observer inclination $I$ 0.82 0.84
magnetic inclination $\iota$ 1.56 1.51
$x_{0}\,(\mathrm{km})$0.51 0.43
$y_{0}\,(\mathrm{km})$0.32 -0.35
$z_{0}\,(\mathrm{km})$ 0.31 0.66
$\log_{10}T_{s}$ 3.09 2.85
Figure 6: (a) The maximum-likelihood hotspot configuration in Case 2: with return current, the color scale indicating the temperature. (b) The maximum-likelihood parameters produced pulse profile of J0030 results comparing with the data
4 Real case study: PSR J0030+0451 and PSR J0437–4715 hotspot map construction
In the sections above, we computed physics-motivated hotspots on the neutron star with an off-center dipole magnetic configuration and derived the temperature distribution of these hotspots. We want to test how well the off-center dipole model can reproduce some of the X-ray observations, and use the results to assess whether higher-order multipole moments are necessary to fully account for the pulsar’s X-ray emission.
In this case study, we focus on two sources: PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437). PSR J0030+0451 was the first millisecond pulsar that NICER source for which detailed pulse profile modeling was performed. The original report on the hotspot configuration of this star, by Riley et al. (2019) and Miller et al. (2019), described one circular hotspot combined with an elongated hotspot, both in the same hemisphere, suggesting the presence of higher-order multipole moments. However, with additional data and new methodologies, Vinciguerra et al. (2024) carried out a follow-up study, and concluded that higher-order multipoles may not be necessary, but a dipole with nontrivial polar cap temperature distribution might be sufficient to explain the data. Whether higher-order multipoles are preferred in the inference method described here for this source is thus of great interest.
The other source, PSR J0437–4715, is the brightest source currently available. Recent analysis of its hotspot configuration by Choudhury et al. (2024) reports a ring-shaped hotspot plus a circular hotspot, which clearly indicates a multipole feature. Investigating whether this shifted dipole field can adequately recover the data for PSR J0437–4715 is also a very intriguing problem.
4.1 Inference methodology
We carry out Bayesian inference using the Markov Chain Monte Carlo (MCMC) method to find the best-fitting parameters for a shifted dipole model that can best reproduce the observed X-ray pulse profiles of the two target sources. To simplify this problem, we fix the mass and radius of these two sources to the maximum likelihood values reported in recent literature. For J0030, these values are $M=1.385M_{\odot}$ and $R=11.70\,\mathrm{km}$, as suggested by the maximum likelihood value from ST-PDT case by Vinciguerra et al. (2024). For J0437, the values are $M=1.418M_{\odot}$ and $R=11.36\,\mathrm{km}$, which corresponded to the result reported by Choudhury et al. (2024). We summarize these fixed model input value in Table 1 for J0030 and J0437 separately. This approach effectively fixes the space-time properties of the pulsar, allowing us to focus solely on the influence of the pulse profile introduced by the hotspot configuration, greatly reducing the parameter space and the amount of computational power required. A comprehensive inference that jointly infers the whole space-time properties and hotspot information will be reserved for future study.
The free parameters need to be explored include the the observer inclination $I$, magnetic inclination angle $\iota$, and the $x$–, $y$–, $z$– direction shifts in the magnetic field: $x_{0}$, $y_{0}$, and $z_{0}$. Since our model predicts only the temperature of the hotspots, we must also account for the temperature of the star’s surface outside the hotspot regions. We assume this surface temperature, $T_{s}$, to be uniform across the whole star surface. This parameter is also free and needs to be explored by our inference.
In principle, the normalization constant of surface temperature, $T_{0}/j_{0}^{1/4}$, should be included as a fitting parameter as well. However, since we only use the normalized 1D pulse profile as our MCMC input, discarding the X-ray spectra, the shape of the light curve only depends on the ratio $T_{0}/T_{s}$. As we have chosen $T_{0}/j_{0}^{1/4}$ to be unity for simplicity, having $T_{s}$ as a free parameter is enough degrees of freedom.
The prior settings for all free parameters are summarized in Table 2. All six free parameters are set with a uniform prior (denoted as $\mathcal{U}$) to reflect the fact that we do not have a clear preference for any specific values. The $x$–, $y$–, $z$– direction shifts have a potential constraint: the total shift of the dipole field should be smaller than the star’s radius. The likelihood of the simulated light curve compared to the real data is expressed as a mean square difference error, reflecting how closely the simulated pulse profile matches the observed data. By combining the prior and likelihood, we can extract the posterior distribution and explore the allowed parameter space.
All the computations of X-ray light curve are based on the X-PSI package (Riley et al., 2023), incorporating a newly developed global temperature map module. This module theoretically handles arbitrary shapes of hotspots and temperature distributions, taking into account effects such as general relativistic light bending and rotational Doppler shift. A similar module has been implemented in Das et al. (2024). The MCMC fitting is performed using the emcee software (Foreman-Mackey et al., 2013).
Figure 7: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shift and $T_{s}$ after applying constraints to the observed pulse profile of J0437. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The dashed line in the 1D corner plots represents the 68% credible interval, and the title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while the temperature $T_{s}$ is in numerical units
Highest Likelihood Case 1 Case 2
observer inclination $I$ 0.64 0.81
magnetic inclination $\iota$ 1.36 1.14
$x_{0}\,(\mathrm{km})$1.21 -1.28
$y_{0}\,(\mathrm{km})$0.36 -0.42
$z_{0}\,(\mathrm{km})$1.32 -0.90
$\log_{10}T_{s}$ 1.89 2.51
Figure 8: (a) Polar cap temperature distribution for the maximum-likelihood hotspot configuration in Case 2: with return current for PSR J0437–4715. (b) The maximum-likelihood parameters produced pulse profile of J0437 results comparing with the data
In these analyses, we set up two cases to explore the effects of the volumetric return current, which is typically neglected in the standard dipole configuration of Lockhart et al. (2019). As discussed above, ignoring this effect leads to a slightly altered hotspot configuration and omits one potential hotspot component. Since its impact has not been proved to be negligible in an off-center dipole scenario, we quantify its influence by performing Bayesian inference under each cases and comparing the resulting maximum-likelihood parameters.
4.2 Inference results: PSR J0030+0451
Figure 5 shows the J0030 full posterior distributions of the model parameters. Whether we perform the fitting without the return current (Case 1) or with return current (Case 2), most parameters converge to pronounced peaks in the posterior space. The observer inclination angle exhibits a consistent peak around $1.0\,\mathrm{rad}$ in both cases, suggesting that this geometry is robust to the presence or absence of return current effects. Interestingly, in Case 2 (with return current), the posterior for the observer inclination is slightly sharper. Conversely, the magnetic inclination angle $\iota$ does not converge to a single dominant peak, showing multiple local maxima. One noteworthy peak lies near $\pi/2$, indicating that placing both hotspots near the equatorial plane can plausibly reproduce the observed light curve. Analysis of the two-dimensional posterior parameter space defined by $I$ and $\iota$ in this corner plot reveals a significant degeneracy between these parameters. Specifically, when $I$ approaches $\pi/2$—indicating an equatorial viewing geometry—the magnetic inclination tends to peak at approximately $0.8\,\mathrm{rad}$. Conversely, as the magnetic inclination is toward $\pi/2$, an alternative peak emerges near $0.6\,\mathrm{rad}$. This behavior underscores the intrinsic degeneracy between the observer’s angle and the magnetic inclination. Moreover, the presence of multiple peaks in the magnetic inclination distribution suggests that its convergence is incomplete, implying that the constraints remain insufficiently precise. Notably, although both the observer angle and the magnetic inclination exhibit prominent features near the $90^{\circ}$ region, these peaks are not directly correlated, further emphasizing the fundamental degeneracy between these two parameters.
For the $x$–, $y$–, and $z$–shifts, the posterior distributions reveal that the star’s dipole displacement is constrained to a relatively small fraction of the stellar radius. However, the $z$–direction shift remains less constrained than the other two, likely because its influence on the pulse profile (see Figure 4) is more moderate. In terms of the surface temperature outside the hotspot, $T_{s}$, both cases yield similar peak values, but the uncertainty range is larger without the return current included (Case 1) than the one with it included (Case 2). This result indicates that including the return current provides tighter constraints on the overall hotspot geometry. Although the more constrained parameter range in Case 2 suggests that including the return current generally provides tighter constraints on the overall hotspot geometry, our maximum-likilhood analysis indicates an magnetic inclination of approximately $90^{\circ}$. At this inclination, the impact of the return current on the hotspot geometry is minimal in this specific regime. Taken together, these findings suggest that the model with return current (Case 2) yields a more better-constrained set of parameters, implying that return current effects may be still essential for accurately modeling the observed emission.
Notably, the highest peaks for the $x$–, $y$–, and $z$– shifts are all very close to zero, indicating that a minimal shift can explain the data quite well. It is also worth commenting on the inclination angle constraint. From gamma-ray observations (Abdo et al., 2013), the magnetic inclination angle presented here is consistent with other observation within 64% uncertainty range. However, given that we are fixing the mass, radius, this agreement is remarkable and interesting.
In the top panel of Figure 6, we present the hotspot configuration corresponding to the maximum-likelihood sample in Case 2, with the detailed model parameters listed in Case 2 column of Figure 5. This configuration shows that both hotspots lie near the equatorial plane, and one appears slightly brighter than the other. The resulting structure is more complex than the configuration inferred by Vinciguerra et al. (2024), which did not include the temperature gradient as incorporated in our approach. This increased complexity leads to a larger magnetic inclination than the previous result, which is approximately $71.7^{\circ}$ inferred from Vinciguerra et al. (2024) ST+PDT maximum likelihood sample.
On the bottom of Figure 6 compares the 1-D pulse profile data with the pulse profiles corresponding to the maximum likelihood samples in with and without return current cases. The bottom panel shows the residuals between the model and the data. The maximum-likelihood model explains the data well, with the main discrepancy occurring slightly in the valley and peak of the pulse profile. Given that we fixed the mass, radius, this is a good result achieved with a minimal number of parameters. There are still some degrees of freedom, such as the temperature normalization constant of each hotspot and the temperature-current relation polynomial index.
This minimal model demonstrates that Case 1 and Case 2 produce similar maximum-likelihood results with comparable residuals. As the magnetic inclination approaches $90^{\circ}$, incorporating a return current produces an almost identical hotspot configuration, which is consistent with our expectations. Overall, the pulse profile of J0030 can potentially be understood within the framework of this off-center dipole field with a good fit, this result also match with the analysis in Vinciguerra et al. (2024). A comprehensive analysis, including a detailed treatment of the joint analysis of the energy spectrum and the background, will be pursued in future studies.
4.3 Inference results: PSR J0437–4715
In the analysis of source J0437, We applied the same inference procedure as for J0030, with the exception of predefined model parameters—namely, mass $M_{\star}$, radius $R_{\star}$, Earth distance $D$, and coordinate frequency $\nu$—which differ from those used in the J0030 analysis. In this study, we keep consistent with the inference set-up for Earth distance and coordinate frequency as described in Choudhury et al. (2024) and adopt their best-model result for mass and radius as our standard inputs, as detailed in Table 1. Figure 7 displays the posterior distributions of the model parameters for the both cases: with and without return current. In both scenarios, the highest likelihood sample is highlighted in the table located at the top-right corner of the figure. The radio band observations provide a precise estimate of the observer inclination angle, approximately $\sim 0.741\,\mathrm{rad}$, which is evident in the tighter constraint of this parameter when the return current is included. Although the inference is not yet fully converged, this preliminary result is encouraging.
For the magnetic inclination angle $\iota$, similar to the J0030 case, the posterior distribution for Case 2 exhibits a pronounced peak near $\pi/2$ along with additional multi-peak features that extend over a broad range, suggesting that this parameter remains poorly constrained by the current observations. The joint two-dimensional posterior distribution for the observer inclination $I$ and magnetic inclination $\iota$ reveals two distinct peaks. One solution shows the magnetic inclination approaching $\pi/2$ with an observer inclination of approximately $0.7\,\mathrm{rad}$, paralleling the behavior seen in the J0030 analysis. The alternative solution, however, features both the observer and magnetic inclinations peaking near $1.0\,\mathrm{rad}$, indicating a nontrivial configuration. Nonetheless, due to the incomplete convergence of the inference for these angles, it remains ambiguous which peak will ultimately dominate the final results. the posterior distributions for the dipole center shifts in the $x$–, $y$–, and $z$–directions are generally broader and the central value are larger compared to the J0030 inference result. This suggests that a larger dipole center shift is required to adequately explain the J0437 data. Furthermore, when comparing the scenarios with and without return current, the inclusion of return current necessitates a larger dipole shift to fit the data while simultaneously providing tighter constraints on the parameter space. This aligns with our findings from the J0030 analysis, indicating a preference for the model with return current due to its more constrained parameter space and better alignment of the maximum-likelihood observer inclination value with radio observations.
All dipole-shift parameters in both cases have significantly decoupled from their priors and exhibit high-peak distributions. Regarding the surface temperature outside the hotspot, $T_{s}$, the posterior distributions are nearly identical between the two scenarios. However, for J0437, it is important to note that our inferred observer inclination is roughly consistent with the prior used in Choudhury et al. (2024), which is informed by radio observations (Reardon et al., 2024). However, if we were to restrict the prior to maintain this consistency, our inference would fail to adequately explain the observational data. There are two possible explanations for this discrepancy: first, the temperature map reported in Choudhury et al. (2024) includes a ring-shaped hotspot accompanying a circular hotspot, introducing distinct multipolar features that an off-center dipole model cannot fully capture. Second, Choudhury et al. (2024) assumes that the radio-observed orbital inclination corresponds directly to the observer inclination angle, based on the assumption that the millisecond pulsar spin axis has aligned with the spin axis of its binary system over long-term evolution. Our results suggest a more flexible assumption, allowing for a potential misalignment.
The resulting hotspot configuration, derived from the highest-likelihood sample, is illustrated in the top of Figure 8 for the case that includes return current. The detailed parameter setup is also shown in the top-right corner of Figure 7 for case 2. This configuration indicates that both the northern and southern hemispheres contain hotspots of similar area, with the northern hotspot having a slightly higher temperature. In addition, each hotspot exhibits an arched component, a feature arising from the inclusion of return current.
Comparing the light curves corresponding to the return-current and no-return-current cases against the observational data reveals that, although both scenarios can produce reasonably convergent posterior distributions, neglecting the return current leads to a noticeably poorer fit—particularly around the half-integer phase. In the no-return-current configuration, an deviation appears at this phase, resulting in larger residuals. This outcome underscores the importance of accounting for the return current in this computation.
In the return-current scenario, the maximum-likelihood sample continues to align reasonably well with the pulse profile data. Together with the posterior distribution in Figure 7, these results suggest that an off-center dipole field configuration, when augmented by return-current effects, can capture the majority of the system’s X-ray emission features.
In all of our analyses, the high computational cost of these inferences has prevented the posterior distributions from fully converging or being rigorously tested across multiple sampler configurations. The main objective of this work is to propose a new physics-motivated off-center-dipole field modeling procedure and demonstrate its feasibility using a state-of-the-art inference pipeline. Consequently, at this proof-of-concept stage, the current inference precision is considered acceptable. Future studies will undertake more advanced, large-scale inference runs that relax the currently fixed mass–radius inputs, in favor of treating the complete set of spacetime-related properties as free parameters to derive comprehensive inference once this physics-motivated framework is fully refined.
5 Discussion and Conlusion
In this study, we employed an off-center dipole magnetic field model to investigate the temperature distribution on the surfaces of neutron stars, focusing on two key sources: PSR J0030+0451 and PSR J0437–4715. Our approach builds on recent development of the theoretical understanding of pulsar magnetospheres and surface heating due to magnetospheric current. Using the approximate force-free dipole solution, our model explicitly computes the temperature distribution of polar caps for an arbitrary off-center magnetic dipole moment of the neutron star, with or without the volumetric return current. The off-center dipole model provides a simplified yet powerful framework for capturing the essential features of the magnetic field configuration, which influences the X-ray emission observed from these stars.
For PSR J0030+0451, our results suggest that the inferred hotspot configuration within the context of an off-center dipole field can yield pulse profiles largely consistent with the observation data. The temperature gradient observed in our model introduces a larger magnetic inclination than that inferred from Fermi gamma-ray data (Abdo et al., 2013; Chen et al., 2020), indicating that higher-order multipole moments might still be necessary to fully account for the star’s multi-wavelength emission.
As for PSR J0437–4715, the inferred hotspot configuration is notably asymmetric, with one hotspot being significantly larger and fainter than the other. This configuration aligns with the observation data, as only one single peak is observed per rotation period. In this particular case, we see that the model with volumetric return current fits the data significantly better, indicating that it may be non-negligible at least in some scenarios.
Our results demonstrate that the off-center dipole field model, despite its simplicity, is already capable of reproducing the observed pulse profiles of both sources with a considerable degree of accuracy. However, for both sources studied in this paper, there are features that are not fully accounted by this simple model. For example, our fitted magnetic inclination for J0030 is somewhat too large, which is in tension with its gamma-ray pulse profile. Our fitted observer inclination for J0437 my not be completely consistent with the very well constrained geometry inferred from the orbital motion. Both cases likely can be further improved with a more rigorous fitting procedure that takes into account the energy-dependent pulse profile. But these inconsistencies may also point to the need for multipole magnetic configurations as those suggested by  Bilous et al. (2019),  Chen et al. (2020) and  Kalapotharakos et al. (2021).
In this study, we directly employed the force-free model developed by Gralla et al. (2017) to compute the magnetospheric current in the off-center dipole scenario. This solution relies on an empirical formula for a crucial quantity $\Lambda$ that determines the magnitude of the current. The empirical formula was found by comparing with first-principles PIC simulations, and its validity in the off-center dipole regime has not been confirmed. Future work should carry out a quantitative comparison between the solution we constructed with global PIC simulations, which may significantly improve the quality of our model.
Similarly, we have employed a non-relativistic definition of the Euler potential $\alpha$ to account for the modifications introduced by the shifted dipole. This assumption ignores the general relativistic correction to the dipole solution, which is difficult to incorporate as the dipole center no longer coincides with the mass center. Nevertheless, we have compared the GR dipole solution from Gralla et al. (2017) with our non-relativistic model, and the deviation is up to around $3\%$. They do not significantly alter the overall distribution of the temperature map or the area of the bright regions. Therefore, we argue that even in the absence of a comprehensive treatment of GR effects, our theoretical computations remain reliable and approximate the hotspot geometry adequately. A more comprehensive GR solution could be pursued in future studies.
Another crucial assumption in this paper is the simple surface heating model where the power is directly proportional to the electric current $j$. Since the polar cap heating is ultimately controlled by the complex QED pair cascade process, it is natural to expect that the temperature profile has more complicated dependence on $j$ than Equation (14). To the best knowledge of the authors, such a phenomenological model has yet to be developed. A detailed microphysical description of how the Poynting flux is dissipated and converted to plasma kinetic energy, which in turn heats the surface of the neutron star, may significantly improve the reliability of future physics-based X-ray pulse profile modeling.
At this stage, conducting a full-scope inference using the energy-resolved 2D pulse profile data with the physics-motivated hotspot model from this paper remains a challenging task due to the extreme computational cost associated with the large dimension of data and computation requirement of the temperature map. Further optimization of the pipeline that computes the energy-resolved pulse profiles from the physical parameters is needed. We defer this nontrivial problem to a future study.
6 Acknowledgements
The authors would like to thanks for the insightful discussions with Anna Watts and Sam Gralla. Special thanks to Pushpita Das, Tuomo Salmi, and Bas Dorsman for their assistance with the usage of the X-PSI software. Additionally, thanks to Serena Vinciguerra for generously providing inference data and guidance on implementing geometric computations. This research is supported by NSF grants DMS-2235457 and AST-2308111, and NASA grant 80NSSC24K1095.

### v3

@@
 In addition to determining the mass and radius of PSR J0030+0451, Riley et al.
+(2019); Miller et al.
 (2019) produced, at the time, the most detailed hotspot map of the neutron star surface derived from observational data.
@@
 A recent investigation by Vinciguerra et al.
-(2024) revisited the hotspot modeling of PSR J0030+0451, finding the hotspots to be consistent with the polar caps of an off-center dipole magnetic field but with a nontrivial temperature distribution.
+(2024) revisited the hotspot modeling of PSR J0030+0451, finding the hotspots to be consistent with the polar caps of an off-center dipole magnetic field but with a nontrivial temperature distribution due to the implementation of new background constraints.
 In particular, the new study found that one of the hotspots has a high-temperature component near its edge.
@@
 Finally, we compute the pulsar’s light curve using the X-PSI global temperature module and incorporate it into pulse-profile modeling process.
-We perform a direct fit to the observed X-ray light-curve data of PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437), aiming to showcase the explanatory power of our off-center dipole temperature map.
+We perform a direct fit to the observed X-ray light-curve data of PSR J0030+0451 (J0030) (Bogdanov et al., 2019) and PSR J0437–4715 (J0437) (Choudhury et al., 2024), aiming to showcase the explanatory power of our off-center dipole temperature map.
 This work illustrates the potential of adopting a physics-based hotspot configuration to model pulsar X-ray data, offering an alternative to the geometric methods currently in use (e.g., those proposed by Riley et al., 2019).
@@
 $J^{\hat{\phi}}$ $=\frac{\Lambda(\alpha,\beta)}{r}\left(\partial_{r}\alpha\partial_{\theta}\beta-\partial_{\theta}\alpha\partial_{r}\beta\right),$ (6)
-Where $(J^{\hat{t}},J^{\hat{r}},J^{\hat{\theta}},J^{\hat{\phi}})$ are the components of the four-current components under the orthonormal basis $(\hat{\bm{t}},\hat{\bm{r}},\hat{\bm{\theta}},\hat{\bm{\phi}})$, and $J^{t}$ is given by:
+[…]
 $\begin{split}J^{t}=&\frac{\Omega-\Omega_{Z}}{r(r-2M)}\left\{\partial_{\theta}\alpha\partial_{\theta}\partial_{\phi}\beta-\partial_{\theta}\beta\partial_{\theta}\partial_{\phi}\alpha\right.\\ &+r(r-2M)\left(\partial_{r}\alpha\partial_{r}\partial_{\phi}\beta-\partial_{r}\beta\partial_{r}\partial_{\phi}\alpha\right)\\ &-\partial_{\phi}\alpha\left[\left(1-\frac{2M}{r}\right)\partial_{r}\left(r^{2}\partial_{r}\beta\right)\right.\\ &\left.+\frac{\partial_{\theta}\left(\sin\theta\partial_{\theta}\beta\right)}{\sin\theta}\right]+\partial_{\phi}\beta\left[\left(1-\frac{2M}{r}\right)\right.\\ &\left.\partial_{r}\left(r^{2}\partial_{r}\alpha\right)+\frac{\partial_{\theta}\left(\sin\theta\partial_{\theta}\alpha\right)}{\sin\theta}\right]\},\end{split}$ (7)
@@
 $M_{\star}$ and $R_{\star}$ are the mass and radius of the neutron star respectively.
+The normalized time component $J^{\hat{t}}$ is the force-free charge density, and can be identified with the Goldreich-Julian charge density $\rho_{\mathrm{GJ}}$ first defined by Goldreich & Julian (1969).
 The function $\Lambda(\alpha,\beta)$ is a conserved quantity along each field line that determines the magnitude of the current density.
@@
 The dipole magnetic moment $\mu$ is defined as $B_{0}\,R_{\star}^{3}$, where $B_{0}$ represents the magnetic field strength on the surface of the neutron star.
-In the subsequent sections, we adopt a value of $B_{0}=5\times 10^{8}\,\mathrm{G}$.
+In the subsequent sections, we adopt a value of $B_{0}=5\times 10^{8}\,\mathrm{G}$ based on the rough spin-down estimation for millisecond pulsar populations.
 Equation (9) is our definition of the pulsar polar cap for the rest of this paper, even when the designated regions are no longer geometrically similar to polar caps in the traditional sense.
@@
 This heating is intimately related to the dissipation of electromagnetic energy, particle acceleration, and the pair production process above the neutron star surface.
-As shown through analytic calculations by of Beloborodov (2008) and simulations by Timokhin & Arons (2013), particle acceleration at the polar cap depends on the ratio $j/c\rho_{\mathrm{GJ}}$, where $j$ is the magnitude of the three-current density and $\rho_{\mathrm{GJ}}$ is the Goldreich-Julian charge density Goldreich & Julian (1969).
+As shown through analytic calculations by of Beloborodov (2008) and simulations by Timokhin & Arons (2013), particle acceleration at the polar cap depends on the ratio $j/c\rho_{\mathrm{GJ}}$, where $j$ is the magnitude of the three-current density defined in Equations (4)–(6), and $\rho_{\mathrm{GJ}}$ is the Goldreich-Julian charge density, which is equivalent to $J^{\hat{t}}$ defined in Equation (3).
 If this ratio is locally greater than 1 (the four-current is spacelike, $J^{2}>0$) or less than 0 ($j$ and $\rho_{\mathrm{GJ}}$ have opposite signs), these regions can build up a large voltage, accelerating particles to very high Lorentz factors, triggering $e^{\pm}$ pair production.
 A fraction of the produced pairs precipate on the polar cap, converting some of the energy dissipated in the gap to heat the neutron star surface.
-On the other hand, if $10<j/c\rho_{\mathrm{GJ}}<$ 1 (the four-current is timelike, $J^{2}<0$), the plasma can flow along the field line with a mildly relativistic Lorentz factor without building up a large voltage (Chen & Beloborodov, 2013).
+On the other hand, if $0<j/c\rho_{\mathrm{GJ}}<1$ (the four-current is timelike, $J^{2}<0$), the plasma can flow along the field line with a mildly relativistic Lorentz factor without building up a large voltage (Chen & Beloborodov, 2013).
 In this paper, we follow the simple prescription used by Lockhart et al.
@@
 etnec
-Figure 2: Pulse profile computed from the centered dipole field in three cases: Uniform hotspot temperature (red), Without return current (blue), With return current (green).
+Figure 2: Pulse profile of bolometric flux computed from the centered dipole field in three cases: Uniform hotspot temperature (red), Without return current (blue), With return current (green).
 All the other parameters to generate this plot are listed in Table 1.
 The magnetic inclination angle $\iota=45^{\circ}$, observer inclination angle $I=90^{\circ}$.
-To compute the X-ray light curves from the surface temperature profiles, we use the open source X-PSI package by Riley et al.
+To compute the X-ray light curves from the surface temperature profiles, we use the open source X-PSI package used by  Riley et al.
 (2019) (https://github.com/xpsi-group/xpsi.git, Riley et al.
-(2023)), which follows a methodology developed over many previous works, including Bogdanov et al.
-(2019); Riley et al.
-(2019); Bogdanov et al.
-(2021); Riley et al.
-(2021).
+(2023)), which follows a methodology developed over many previous works (Riley, 2019).
+This code was bechmarked in Bogdanov et al.
+(2019, 2021), and used in analysis Riley et al.
+(2019, 2021).
 To describe the shape of the hotspots, the X-PSI package normally utilizes a model that represents the hot-spot surface through overlapping spherical caps with uniform temperature, but it also has the capability to accommodate irregular shapes of hot-spots by specifying the temperature distribution over the whole stellar surface  (see e.g.
@@
 Figure 3: 4-current density squared $J^{2}$ and temperature $T$ distribution of shifted hotspot projected on the star surface, with magnetic inclination $\iota=45^{\circ}$ and dipole shift $\delta=R_{\star}/2$ along different directions.
-Panels (a) and (b): $x$–direction shift; panels (c) and (d): $y$–direction; panels (e) and (f): $z$–direction shift.
+Panels (a) and (b): xx–direction shift; panels (c) and (d): yy–direction; panels (e) and (f): zz–direction shift.
 Each color map is normalized by the maximum value of $J^{2}$ and $T$.
-Figure 4: Pulse profile computed from the shifted dipole field in three cases: $x$–direction shift (blue), $y$–direction shift (red) and $z$–direction shift (green), comparing with centered dipole field computation without including any shift (black dashed).
+Figure 4: Pulse profile computed from the shifted dipole field in three cases: xx–direction shift (blue), yy–direction shift (red) and zz–direction shift (green), comparing with centered dipole field computation without including any shift (black dashed).
 All the other parameters to generate this plot are listed in Table 1 with magnetic inclination $\iota=45^{\circ}$, observer inclination $I=90^{\circ}$
@@
 Figure 3 illustrates the effect of a shift $\delta=R_{\star}/2$ along the $x$, $y$, and $z$ directions on the polar cap 4-current density structure and temperature distribution.
-For example, in the panel (a), we present the $x$–direction shift field 4-current density distribution across the polar cap on the stellar surface.
+For example, in the panel (a), we present the xx–direction shift field 4-current density distribution across the polar cap on the stellar surface.
 Compared to the centered dipole configuration, the north polar cap is greatly extended to a larger area, with a single hotspot at its center.
@@
 The panel (b) shows that the south polar cap , despite being much smaller, has a significantly higher temperature than the north.
-This is consistent with the expectation that regions located closer to the dipole center experience stronger magnetic fields and thus have stronger heating.
-Similar asymmetry between the temperatures of the two polar caps can be seen in the $z$–shifted dipole case as well (see panels (e) and (f) of Figure 3).
-The four-current density and temperature distribution for the $y$–shift scenario $\delta_{y}=R_{\star}/2$ in a dipole field are shown in the panels (c) and (d) of Figure 3.
+This follows from the fact that, in our model, the local surface temperature is set by the magnitude of the current density $j$.
+Each open magnetic flux surface carries a conserved current $\Lambda(\alpha,\beta)$, fixed in the far zone where the field is symmetric across the current sheet.
+Because both caps intercept the same total current, squeezing one cap into a smaller area inevitably raises $j$ there, and thus (after the red-shift factor $\sqrt{1-2M/R_{\star}}$ is applied) delivers more heating per unit area.
+Similar asymmetry between the temperatures of the two polar caps can be seen in the zz–shifted dipole case as well (see panels (e) and (f) of Figure 3).
+The four-current density and temperature distribution for the yy–shift scenario $\delta_{y}=R_{\star}/2$ in a dipole field are shown in the panels (c) and (d) of Figure 3.
 The two polar caps remain symmetric, and generally similar to the unshifted case shown in Figure 1.
@@
 This displacement changes the distance between the two peaks in the X-ray light curve in one rotation period, as can be seen in Figure 4.
-Such a distinct outcome for a $y$–direction shift arises naturally because the star’s dipole moment lies in the $x-z$ plane, so shifting perpendicular to that plane yields behavior not seen in within-plane shifts.
-The panels (e) and (f) of Figure 3 show how a shift $\delta_{z}=R_{\star}/2$ along the $z$–direction affects the hotspot current density distribution and temperature map.
-In panel (e), the two polar caps display noticeably different areas, with the north polar cap undergoing an enlargement similar to what was observed in the $x$–shift scenario, and southern polar cap also exhibited a larger current density similar with what we see in $x$–direction shift.
-The panel (f) reveals that, in this configuration, the south hotspot appears significantly brighter, whereas the north hotspot—despite having a larger area— is comparatively fainter, similiar to the features seen in the $x$–shift case.
-The X-ray pulse profiles for shifts in the $x$–, $y$–, and $z$–directions are computed and presented in Figure 4 for a neutron star with magnetic inclination angle $\iota=45^{\circ}$, as seen by an observer at angle $I=90^{\circ}$ with respect to the rotation axis.
+Such a distinct outcome for a yy–direction shift arises naturally because the star’s dipole moment lies in the xzx-z plane, so shifting perpendicular to that plane yields behavior not seen in within-plane shifts.
+The panels (e) and (f) of Figure 3 show how a shift $\delta_{z}=R_{\star}/2$ along the zz–direction affects the hotspot current density distribution and temperature map.
+In panel (e), the two polar caps display noticeably different areas, with the north polar cap undergoing an enlargement similar to what was observed in the xx–shift scenario, and southern polar cap also exhibited a larger current density similar with what we see in xx–direction shift.
+The panel (f) reveals that, in this configuration, the south hotspot appears significantly brighter, whereas the north hotspot—despite having a larger area— is comparatively fainter, similiar to the features seen in the xx–shift case.
+The X-ray pulse profiles for shifts in the xx–, yy–, and zz–directions are computed and presented in Figure 4 for a neutron star with magnetic inclination angle $\iota=45^{\circ}$, as seen by an observer at angle $I=90^{\circ}$ with respect to the rotation axis.
 Additional parameters, such as the mass and radius of the neutron star, are required to determine its external spacetime, thereby allowing us to generate the pulse profiles using the X-PSI package.
 We used the parameters of the pulsar J0030+0451 for these light curves, and the values are summarized in Table 1.
-Without any normalization added, our model indicates that for an $x$–direction shift, X-ray emission from the south polar cap dominates, yielding a significantly larger peak-to-valley ratio than the centered-dipole light curve (shown as the dashed curve).
+Without any normalization added, our model indicates that for an xx–direction shift, X-ray emission from the south polar cap dominates, yielding a significantly larger peak-to-valley ratio than the centered-dipole light curve (shown as the dashed curve).
 Although the south polar cap is smaller in area, it is brighter due to the locally enhanced magnetic field.
-Consequently, the maximum peak in the $x$–shift configuration is nearly twice that of the centered-dipole setup.
-In the $y$–direction shift scenario, the north and south polar caps move closer, merging into a single, highly boosted emission peak; Figure 4 shows that this configuration produces both the highest peak and the largest peak-to-valley ratio, implying a higher total flux than in the other shift cases.
-Finally, the $z$–direction shift presents an interesting contrast.
+Consequently, the maximum peak in the xx–shift configuration is nearly twice that of the centered-dipole setup.
+In the yy–direction shift scenario, the north and south polar caps move closer, merging into a single, highly boosted emission peak; Figure 4 shows that this configuration produces both the highest peak and the largest peak-to-valley ratio, implying a higher total flux than in the other shift cases.
+Finally, the zz–direction shift presents an interesting contrast.
 Although the north polar cap appears fainter than the south cap in Figure 3, the observer inclination $I=90^{\circ}$ (equatorial viewing) brings the north cap closer to the line of sight.
@@
 $B$ in G [Gauss], is the surface magnetic field that take into the computation of magnetic moment.
+The frequency and earth distance of J0030 are derived from radio observation Arzoumanian et al.
+(2018), and the frequency and earth distance of J0437 was from Reardon et al.
+(2024).
   Model Parameters   Prior
-      observer inclination $I$       $\mathcal{U}(0,\frac{\pi}{2})$
+  observer inclination II   $\mathcal{U}(0,\frac{\pi}{2})$
   magnetic inclination $\iota$   $\mathcal{U}(0,\frac{\pi}{2})$
@@
   $\log_{10}T_{s}$   $\mathcal{U}(0,10)$
-Table 2: Summary of Inference Model Parameters and Prior Settings: The observer inclination is denoted by $I$ and the magnetic inclination by $\iota$.
-The off-center dipole displacements along the $x$–, $y$–, and $z$–axes are represented by $x_{0}$, $y_{0}$, and $z_{0}$, respectively.
+Table 2: Summary of Inference Model Parameters and Prior Settings: The observer inclination is denoted by II and the magnetic inclination by $\iota$.
+The off-center dipole displacements along the xx–, yy–, and zz–axes are represented by $x_{0}$, $y_{0}$, and $z_{0}$, respectively.
 $T_{s}$ denotes the temperature of the neutron star’s non-hotspot region (in arbitrary units).
 Here, $\mathcal{U}$ denotes a uniform distribution.
-Figure 5: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shifts and $\log_{10}T_{s}$ after applying constraints to the observed pulse profile of J0030.
+Figure 5: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), xxyyzz direction shifts and $\log_{10}T_{s}$ after applying constraints to the observed pulse profile of J0030.
 Blue is the posterior with Case 1: without return current.
@@
 The title of this plot indicates the median of the distribution as well as the range of the 68% credible interval.
-Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while $T_{s}$ is in numerical units.
-Highest Likelihood Case 1 Case 2
-observer inclination $I$ 0.82 0.84
-magnetic inclination $\iota$ 1.56 1.51
-$x_{0}\,(\mathrm{km})$0.51 0.43
-$y_{0}\,(\mathrm{km})$0.32 -0.35
-$z_{0}\,(\mathrm{km})$ 0.31 0.66
-$\log_{10}T_{s}$ 3.09 2.85
+Here, xxyyzz shifts are measured in $\mathrm{km}$, while $T_{s}$ is in numerical units.
+Maximum Likelihood Case 1 Case 2
+observer inclination II 0.78 0.68
+magnetic inclination $\iota$ 1.52 0.98
+$x_{0}\,(\mathrm{km})$ -0.35 0.66
+$y_{0}\,(\mathrm{km})$ -0.22 -0.29
+$z_{0}\,(\mathrm{km})$ -0.25 -0.16
+$\log_{10}T_{s}$ 2.40 2.62
 Figure 6: (a) The maximum-likelihood hotspot configuration in Case 2: with return current, the color scale indicating the temperature.
@@
 (2019), described one circular hotspot combined with an elongated hotspot, both in the same hemisphere, suggesting the presence of higher-order multipole moments.
-However, with additional data and new methodologies, Vinciguerra et al.
+However, with new background constraint from XMM-Newton data, Vinciguerra et al.
 (2024) carried out a follow-up study, and concluded that higher-order multipoles may not be necessary, but a dipole with nontrivial polar cap temperature distribution might be sufficient to explain the data.
@@
 This approach effectively fixes the space-time properties of the pulsar, allowing us to focus solely on the influence of the pulse profile introduced by the hotspot configuration, greatly reducing the parameter space and the amount of computational power required.
-A comprehensive inference that jointly infers the whole space-time properties and hotspot information will be reserved for future study.
-The free parameters need to be explored include the the observer inclination $I$, magnetic inclination angle $\iota$, and the $x$–, $y$–, $z$– direction shifts in the magnetic field: $x_{0}$, $y_{0}$, and $z_{0}$.
+The MCMC chains presented below are exploratory just sufficient to demonstrate that the physics-motivated / shifted-dipole model presented here can be sampled with a standard Bayesian pipeline.
+We have not yet established statistical convergence across alternative sampling schemes, even though the chains display qualitative convergence indicators, and Figures 7 still show minimal visible sampling artifacts despite cosmetic changes (wider axes and KDE contours) adopted to minimize distraction.
+A full-scale analysis with longer chains, multiple sampler configurations, and the mass–radius parameters set free is deferred to future work.
+The free parameters need to be explored include the the observer inclination $I$, magnetic inclination angle $\iota$, and the xx–, yy–, zz– direction shifts in the magnetic field: $x_{0}$, $y_{0}$, and $z_{0}$.
 Since our model predicts only the temperature of the hotspots, we must also account for the temperature of the star’s surface outside the hotspot regions.
-We assume this surface temperature, $T_{s}$, to be uniform across the whole star surface.
+We assume this surface temperature, $T_{s}$, to be uniform and finite across the whole star surface.
 This parameter is also free and needs to be explored by our inference.
@@
 All six free parameters are set with a uniform prior (denoted as $\mathcal{U}$) to reflect the fact that we do not have a clear preference for any specific values.
-The $x$–, $y$–, $z$– direction shifts have a potential constraint: the total shift of the dipole field should be smaller than the star’s radius.
+The xx–, yy–, zz– direction shifts have a potential constraint: the total shift of the dipole field should be smaller than the star’s radius.
 The likelihood of the simulated light curve compared to the real data is expressed as a mean square difference error, reflecting how closely the simulated pulse profile matches the observed data.
@@
 This module theoretically handles arbitrary shapes of hotspots and temperature distributions, taking into account effects such as general relativistic light bending and rotational Doppler shift.
+Here we used blackbody as the emission model without including realistic atmosphere model to reduce the computation time.
 A similar module has been implemented in Das et al.
@@
 The MCMC fitting is performed using the emcee software (Foreman-Mackey et al., 2013).
-Figure 7: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shift and $T_{s}$ after applying constraints to the observed pulse profile of J0437.
+Figure 7: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), xxyyzz direction shift and $T_{s}$ after applying constraints to the observed pulse profile of J0437.
 Blue is the posterior with Case 1: without return current.
@@
 The dashed line in the 1D corner plots represents the 68% credible interval, and the title of this plot indicates the median of the distribution as well as the range of the 68% credible interval.
-Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while the temperature $T_{s}$ is in numerical units
-Highest Likelihood Case 1 Case 2
-observer inclination $I$ 0.64 0.81
-magnetic inclination $\iota$ 1.36 1.14
-$x_{0}\,(\mathrm{km})$1.21 -1.28
-$y_{0}\,(\mathrm{km})$0.36 -0.42
-$z_{0}\,(\mathrm{km})$1.32 -0.90
-$\log_{10}T_{s}$ 1.89 2.51
+Here, xxyyzz shifts are measured in $\mathrm{km}$, while the temperature $T_{s}$ is in numerical units
+Maximum Likelihood Case 1 Case 2
+observer inclination II 0.40 0.59
+magnetic inclination $\iota$ 1.22 1.09
+$x_{0}\,(\mathrm{km})$ 1.90 -1.85
+$y_{0}\,(\mathrm{km})$ -0.67 -2.07
+$z_{0}\,(\mathrm{km})$ -0.27 -0.29
+$\log_{10}T_{s}$ 2.33 2.53
 Figure 8: (a) Polar cap temperature distribution for the maximum-likelihood hotspot configuration in Case 2: with return current for PSR J0437–4715.
-(b) The maximum-likelihood parameters produced pulse profile of J0437 results comparing with the data
+(b) The maximum-likelihood parameters produced pulse profile of J0437 results comparing with the bolometric flux data in NICER band.
+A companion animated figure showing the evolution of this hotspot configuration as the off‑center distance increases toward the best‑fit state is available online; an archived copy is hosted on Zenodo: https://doi.org/10.5281/zenodo.16548609 (catalog doi:10.5281/zenodo.16548609).
 In these analyses, we set up two cases to explore the effects of the volumetric return current, which is typically neglected in the standard dipole configuration of Lockhart et al.
@@
 The observer inclination angle exhibits a consistent peak around $1.0\,\mathrm{rad}$ in both cases, suggesting that this geometry is robust to the presence or absence of return current effects.
-Interestingly, in Case 2 (with return current), the posterior for the observer inclination is slightly sharper.
-Conversely, the magnetic inclination angle $\iota$ does not converge to a single dominant peak, showing multiple local maxima.
-One noteworthy peak lies near $\pi/2$, indicating that placing both hotspots near the equatorial plane can plausibly reproduce the observed light curve.
-Analysis of the two-dimensional posterior parameter space defined by $I$ and $\iota$ in this corner plot reveals a significant degeneracy between these parameters.
-Specifically, when $I$ approaches $\pi/2$—indicating an equatorial viewing geometry—the magnetic inclination tends to peak at approximately $0.8\,\mathrm{rad}$.
-Conversely, as the magnetic inclination is toward $\pi/2$, an alternative peak emerges near $0.6\,\mathrm{rad}$.
-This behavior underscores the intrinsic degeneracy between the observer’s angle and the magnetic inclination.
-Moreover, the presence of multiple peaks in the magnetic inclination distribution suggests that its convergence is incomplete, implying that the constraints remain insufficiently precise.
-Notably, although both the observer angle and the magnetic inclination exhibit prominent features near the $90^{\circ}$ region, these peaks are not directly correlated, further emphasizing the fundamental degeneracy between these two parameters.
-For the $x$–, $y$–, and $z$–shifts, the posterior distributions reveal that the star’s dipole displacement is constrained to a relatively small fraction of the stellar radius.
-However, the $z$–direction shift remains less constrained than the other two, likely because its influence on the pulse profile (see Figure 4) is more moderate.
+The posterior distribution of the magnetic inclination angle, $\iota$, converges to a single dominant peak near $1.0$, indicating satisfactory convergence even without repeated runs under alternative sampling configurations.
+For the $x$-, $y$-, and $z$-shift parameters, the posterior distributions indicate that, for J0030+0451, the stellar dipole is displaced by only a small fraction of the stellar radius: all shifts are well constrained within 1 km (i.e., $10\%$ of $R_{\star}$).
 In terms of the surface temperature outside the hotspot, $T_{s}$, both cases yield similar peak values, but the uncertainty range is larger without the return current included (Case 1) than the one with it included (Case 2).
-This result indicates that including the return current provides tighter constraints on the overall hotspot geometry.
-Although the more constrained parameter range in Case 2 suggests that including the return current generally provides tighter constraints on the overall hotspot geometry, our maximum-likilhood analysis indicates an magnetic inclination of approximately $90^{\circ}$.
-At this inclination, the impact of the return current on the hotspot geometry is minimal in this specific regime.
-Taken together, these findings suggest that the model with return current (Case 2) yields a more better-constrained set of parameters, implying that return current effects may be still essential for accurately modeling the observed emission.
-Notably, the highest peaks for the $x$–, $y$–, and $z$– shifts are all very close to zero, indicating that a minimal shift can explain the data quite well.
+This result indicates that including the return current provides tighter constraints on the overall star surface background temperature.
+Our maximum-likelihood analysis favors a magnetic inclination in which the model without a return current yields $\iota\approx 90^{\circ}$, whereas the model including a return current gives a more plausible value of $\iota\sim 56^{\circ}$.
+This contrast strongly supports the necessity of incorporating return-current physics when using first-principles, physics-motivated hotspot models to interpret the observations.
+Taken together, these findings suggest that the model with return current (Case 2) yields a more reasonable set of parameters, implying that return current effects may be essential for accurately modeling the observed emission.
+Notably, the highest peaks for the xx–, yy–, and zz– shifts are all very close to zero, indicating that a minimal shift can explain the data quite well.
 It is also worth commenting on the inclination angle constraint.
-From gamma-ray observations (Abdo et al., 2013), the magnetic inclination angle presented here is consistent with other observation within 64% uncertainty range.
-However, given that we are fixing the mass, radius, this agreement is remarkable and interesting.
+From gamma-ray observations (Abdo et al., 2013), the magnetic inclination angle presented here is smaller other observation within 64% uncertainty range.
+However, given that we are fixing the mass, radius, this is still remarkable and interesting.
 In the top panel of Figure 6, we present the hotspot configuration corresponding to the maximum-likelihood sample in Case 2, with the detailed model parameters listed in Case 2 column of Figure 5.
-This configuration shows that both hotspots lie near the equatorial plane, and one appears slightly brighter than the other.
+This configuration shows that both hotspots lie near the equatorial plane, and one appears significantly brighter but smaller than the other.
 The resulting structure is more complex than the configuration inferred by Vinciguerra et al.
 (2024), which did not include the temperature gradient as incorporated in our approach.
-This increased complexity leads to a larger magnetic inclination than the previous result, which is approximately $71.7^{\circ}$ inferred from Vinciguerra et al.
+This increased complexity leads to a smaller magnetic inclination than the previous result, which is approximately $71.7^{\circ}$ inferred from Vinciguerra et al.
 (2024) ST+PDT maximum likelihood sample.
@@
 This minimal model demonstrates that Case 1 and Case 2 produce similar maximum-likelihood results with comparable residuals.
-As the magnetic inclination approaches $90^{\circ}$, incorporating a return current produces an almost identical hotspot configuration, which is consistent with our expectations.
 Overall, the pulse profile of J0030 can potentially be understood within the framework of this off-center dipole field with a good fit, this result also match with the analysis in Vinciguerra et al.
@@
 Although the inference is not yet fully converged, this preliminary result is encouraging.
-For the magnetic inclination angle $\iota$, similar to the J0030 case, the posterior distribution for Case 2 exhibits a pronounced peak near $\pi/2$ along with additional multi-peak features that extend over a broad range, suggesting that this parameter remains poorly constrained by the current observations.
-The joint two-dimensional posterior distribution for the observer inclination $I$ and magnetic inclination $\iota$ reveals two distinct peaks.
-One solution shows the magnetic inclination approaching $\pi/2$ with an observer inclination of approximately $0.7\,\mathrm{rad}$, paralleling the behavior seen in the J0030 analysis.
-The alternative solution, however, features both the observer and magnetic inclinations peaking near $1.0\,\mathrm{rad}$, indicating a nontrivial configuration.
-Nonetheless, due to the incomplete convergence of the inference for these angles, it remains ambiguous which peak will ultimately dominate the final results.
-the posterior distributions for the dipole center shifts in the $x$–, $y$–, and $z$–directions are generally broader and the central value are larger compared to the J0030 inference result.
+For the magnetic inclination angle $\iota$, similar to the J0030 case, the posterior distribution for Case 2 exhibits a pronounced peak near $1.0$ with a wide peak, suggesting that this parameter remains less constrained by the current observations.
+The 2D posterior distribution features both the observer and magnetic inclinations peaking near $1.0\,\mathrm{rad}$, indicating a nontrivial configuration.
+The posterior distributions for the dipole center shifts in the xx–, yy–, and zz–directions are generally broader and the central value are larger compared to the J0030 inference result.
 This suggests that a larger dipole center shift is required to adequately explain the J0437 data.
-Furthermore, when comparing the scenarios with and without return current, the inclusion of return current necessitates a larger dipole shift to fit the data while simultaneously providing tighter constraints on the parameter space.
+Furthermore, in direct comparisons between models with and without a return current, the inclusion of a return current can help the dipole-shift posterior contract to a predominant peak and simultaneously yields tighter constraints across the parameter space, whereas, in the case without a return current, the posterior is far more dispersed and exhibits poor convergence, further highlighting the importance of including a return current.
+Collectively, these results indicate improved identifiability of the parameters when return-current physics is incorporated.
 This aligns with our findings from the J0030 analysis, indicating a preference for the model with return current due to its more constrained parameter space and better alignment of the maximum-likelihood observer inclination value with radio observations.
-All dipole-shift parameters in both cases have significantly decoupled from their priors and exhibit high-peak distributions.
-Regarding the surface temperature outside the hotspot, $T_{s}$, the posterior distributions are nearly identical between the two scenarios.
+All dipole-shift parameters in both models are strongly decoupled from their priors and exhibit sharply peaked posteriors.
+For the surface temperature outside the hotspot, $T_{s}$, the posterior distributions are nearly identical across the two scenarios, however, in the model without a return current the distribution is noisier, with some samples extending to very large $T_{s}$ and a small secondary peak near $T_{s}\sim 4$.
 However, for J0437, it is important to note that our inferred observer inclination is roughly consistent with the prior used in Choudhury et al.
@@
 The detailed parameter setup is also shown in the top-right corner of Figure 7 for case 2.
-This configuration indicates that both the northern and southern hemispheres contain hotspots of similar area, with the northern hotspot having a slightly higher temperature.
-In addition, each hotspot exhibits an arched component, a feature arising from the inclusion of return current.
+This configuration indicates that the northern hotspot exhibits a ring-shaped bright pattern, whereas the southern hotspot spans a larger area but is substantially fainter.
+Moreover, an additional asymmetry appears as a small emission region in the southeastern quadrant.
+This morphology is particularly intriguing, since the best-fitting result of Choudhury et al.
+(2024) likewise favors a configuration that is qualitatively similar and features an asymmetric hotspot pattern.
+To better demonstrate how an off-centered dipole field yields the observed asymmetric hotspot geometry, we provide a companion video that visualizes the evolution of the best-fitting hotspot configuration as the off-center distance is increased to its optimal value.
+The video is available on Zenodo: https://doi.org/10.5281/zenodo.16548609 (catalog doi:10.5281/zenodo.16548609).
+In addition, each hotspot exhibits an important arched component, a feature only arising from the inclusion of return current.
 Comparing the light curves corresponding to the return-current and no-return-current cases against the observational data reveals that, although both scenarios can produce reasonably convergent posterior distributions, neglecting the return current leads to a noticeably poorer fit—particularly around the half-integer phase.
@@
 Together with the posterior distribution in Figure 7, these results suggest that an off-center dipole field configuration, when augmented by return-current effects, can capture the majority of the system’s X-ray emission features.
-In all of our analyses, the high computational cost of these inferences has prevented the posterior distributions from fully converging or being rigorously tested across multiple sampler configurations.
+In all of our analyses, the high computational cost of these inferences has prevented being rigorously tested across multiple sampler configurations.
 The main objective of this work is to propose a new physics-motivated off-center-dipole field modeling procedure and demonstrate its feasibility using a state-of-the-art inference pipeline.
-Consequently, at this proof-of-concept stage, the current inference precision is considered acceptable.
+Consequently, at this proof‑of‑concept stage, the parameter estimates should be regarded as illustrative rather than definitive: the chains achieve a plausibly reasonable precision and display apparent convergence, yet a rigorous, multi sampling setup convergence demonstration is deferred to future work.
 Future studies will undertake more advanced, large-scale inference runs that relax the currently fixed mass–radius inputs, in favor of treating the complete set of spacetime-related properties as free parameters to derive comprehensive inference once this physics-motivated framework is fully refined.
@@
 For PSR J0030+0451, our results suggest that the inferred hotspot configuration within the context of an off-center dipole field can yield pulse profiles largely consistent with the observation data.
-The temperature gradient observed in our model introduces a larger magnetic inclination than that inferred from Fermi gamma-ray data (Abdo et al., 2013; Chen et al., 2020), indicating that higher-order multipole moments might still be necessary to fully account for the star’s multi-wavelength emission.
-As for PSR J0437–4715, the inferred hotspot configuration is notably asymmetric, with one hotspot being significantly larger and fainter than the other.
-This configuration aligns with the observation data, as only one single peak is observed per rotation period.
+The temperature gradient observed in our model introduces a smaller magnetic inclination than that inferred from Fermi gamma-ray data (Abdo et al., 2013; Chen et al., 2020), indicating that higher-order multipole moments might still be necessary to fully account for the star’s multi-wavelength emission.
+As for PSR J0437–4715, the inferred hotspot configuration is notably asymmetric, with one hotspot exhibiting an interesting asymmetric structure that includes a small component located in the southeastern quadrant of the hotspot.
+This configuration surprisingly aligns with the best-fit hotspot configuration reported by Choudhury et al.
+(2024).
 In this particular case, we see that the model with volumetric return current fits the data significantly better, indicating that it may be non-negligible at least in some scenarios.
@@
 However, for both sources studied in this paper, there are features that are not fully accounted by this simple model.
-For example, our fitted magnetic inclination for J0030 is somewhat too large, which is in tension with its gamma-ray pulse profile.
-Our fitted observer inclination for J0437 my not be completely consistent with the very well constrained geometry inferred from the orbital motion.
+For example, our fitted magnetic inclination for J0030 is somewhat small, which is in tension with its gamma-ray pulse profile.
+Our fitted observer inclination for J0437 not be completely consistent with the very well constrained geometry inferred from the orbital motion.
 Both cases likely can be further improved with a more rigorous fitting procedure that takes into account the energy-dependent pulse profile.
@@
 The authors would like to thanks for the insightful discussions with Anna Watts and Sam Gralla.
+We also thank the anonymous referee for their comments that helped improve the manuscript.
 Special thanks to Pushpita Das, Tuomo Salmi, and Bas Dorsman for their assistance with the usage of the X-PSI software.